# Universal and Experiment-calibrated Prediction of XANES through Crystal Graph Neural Network and Transfer Learning Strategy

### v1

@@ Abstract
Theoretical simulation is helpful for accurate interpretation of experimental X-ray absorption near-edge structure (XANES) spectra that contain rich atomic and electronic structure information of materials. However, current simulation methods are usually too complex to give the needed accuracy and timeliness when a large amount of data need to be analyzed, such as for in-situ characterization of battery materials. To address these problems, artificial intelligence (AI) models have been developed for XANES prediction. However, instead of using experimental XANES data, the existing models are trained using simulated data, resulting in significant discrepancies between the predicted and experimental spectra. Also, the universality across different elements has not been well studied for such models. In this work, we firstly establish a crystal graph neural network, pre-trained on simulated XANES data covering 48 elements, to achieve universal XANES prediction with a low average relative square error of 0.020223; and then utilize transfer learning to calibrate the model using a small experimental XANES dataset. After calibration, the edge energy misalignment error of the predicted S, Ti and Fe K edge XANES is significantly reduced by about 55%. The method demonstrated in this work opens up a new way to achieve fast, universal, and experiment-calibrated XANES prediction.
@@ I Introduction
X-ray absorption spectroscopy (XAS) is a powerful tool to detect the electronic structures and atomic geometry of materials, including atom valence, bond length and coordination number[RN150]. It has the advantages of non-destructive, element selectivity, and high sensitivity[RN198]. In addition, XAS is available for various sample types, such as bulk materials, powders, and solutions[RN315]. Therefore, it has been widely applied to study the problems in different areas, for example, low dimensional materials characterization[RN317, RN318], catalytic mechanism analysis[RN244, RN195], relics identification[RN316], and biomedical analyses[RN241]. Especially, these unique advantages enable XAS to combine with the operando/in-situ techniques for real-time characterization of some complex systems, like electrochemistry reactions in batteries[RN134, RN133, RN174, RN358]. According to the energy region with respect of the absorption edge, XAS can be divided into X-ray absorption near-edge structure (XANES) and extended X-ray absorption spectroscopy (EXAFS). EXAFS is mainly dominated by single scattering events, where coordination numbers and interatomic distances can be derived from Fourier transform for some materials with periodic structure[RN176, RN124]. XANES is related to strong scattering processes as well as local atomic resonances, which is more complicated than EXAFS and hard to directly transformed to structure parameters[RN230]. Commonly, interpretation of experimental XANES is based on some empirical features, such as comparing edge shifts and peak positions with referenced spectra of standard materials[RN176].
@@
With the development of XAS theory and the corresponding calculation software like FEFF9[RN217], FDMNES[Bunău_2009], and OCEAN[RN251], the XANES can be quantitively simulated from a given crystal or molecule structure and taken as the referenced spectra in analysis. It helps to gain deeper insight into the relationship between local environment of absorbing atoms with spectra features. However, such calculation is time-consuming when the materials structure is complex. And the calculated spectra are sensitive to the selected approximation methods and calculation parameters, where a tedious parameters optimization process is needed to obtain reasonable results. In addition, XANES simulation faces challenges in aligning absolute energy scales and accounting for experimental broadening effects, resulting in inevitable discrepancies compared with experiments[RN249]. Without careful experimental calibration, the simulated XANES may yield false edge and peak features, misleading the XANES analysis. These shortcomings of current simulation methods have brought significant challenges for efficient and accurate XANES interpretation. This especially limits the in-situ XANES study of some essential but complex systems requiring massive data analysis, such as battery materials that experience constant structure changes during charge and discharge processes.
@@
The development of artificial intelligent (AI) has opened up a new paradigm to solve the similar complex problems in different areas, such as AlphaFold for protein structure prediction[RN253], GNoME for material stable structures discovery[RN256], and DeepH for Hamiltonian calculation in DFT[RN255]. Researchers have also made some attempts to apply AI methods in XANES prediction to avoid the complex calculation process. These developed models cover the predictions on C K edge XANES of amorphous carbon and organic molecules[RN121, RN142], O and N K edge XANES of molecules[RN138], and 3d transition metals K edge XANES of transition metal compounds[RN365, RN192, RN367]. Although those models offer a more efficient and convenient way to obtain XANES, their universality is challenged for they only focus on limited range of absorption elements. The universality of AI based XANES prediction on other common elements need to be further studied. In addition, due to the shortage of experimental spectra, the existing models are trained based on the high throughput simulated XANES, where the spectra features and energy scale are not accurately calibrated. The uncorrected discrepancies between the training data and experiment measurement challenge the reliability of existing XANES prediction models. Therefore, the universality and fidelity of AI based XANES prediction should be improved before it can be put into practical use.
@@
In this work, we explored the universal and experiment-calibrated prediction of XANES through crystal graph neural network and transfer learning strategy. A crystal graph neural network model for XANES prediction (CGXAS) is firstly established, and the universality of CGXAS on common detected elements of K edge XANES is studied. Then, transfer learning strategy is adopted to calibrate the predicted XANES using a small experimental dataset. The framework of CGXAS and the training flowchart is represented in Fig.1. In CGXAS, the graph neural network (GNN) encodes the local environment information into a latent site vector for each atom in the crystal graph. Then, the multilayer perceptron (MLP) decodes the site vector into the corresponded K edge XANES of each site. The successive mask pooling layer is used to average the XANES at each site according to the element, enabling XANES prediction for different elements in one network. A universal dataset with 341405 simulated XANES covering 48 elements[RN219] is used for CGXAS training and evaluation. This study compares three distinct approaches for achieving universal XANES prediction: (1) Training a single model (CGXAS_Uni) on the full dataset for universal XANES prediction; (2) Training individual element-specific CGXAS models using the element-specific subsets and assembling them for universal XANES prediction, where model for element X is denoted as CGXAS_Spec_X; (3) Fine-tuning CGXAS_Uni model on each element-specific subset, leading to CGXAS_FT_X model that represents the fine-tuned version for element X. Among the three models, the single CGXAS_Uni model demonstrated the best overall accuracy and universality. To study the interpretability of CGXAS, the t-SNE method is utilized to analyze the latent site vectors. Furthermore, to bridge the gap between simulation and experiment, transfer learning is also adopted to adapt the pretrained CGXAS_Uni model with a small experimental XANES dataset of element X, where the model after transfer learning is denoted as CGXAS_Exp_X. Compared with traditional learning process, transfer learning can transfer the knowledge learned in other domains to enhance the prediction performance on target domain[RN319]. Since the prediction performance of neural network is closely related with training data size, transferring simulated XANES knowledge to experimental XANES prediction can compensate the insufficiency in experimental data quantity. Here, we demonstrate the ability of transfer learning to calibrate the XANES prediction of S, Ti and Fe K edge XANES using only 48, 40 and 45 experiment spectra for each element, where S, Ti, and Fe are the critical elements in the study of energy and catalysis. The calibrated CGXAS_Exp_S, CGXAS_Exp_Ti, and CGXAS_Exp_Fe models reduce the edge energy misalignment error of the predicted XANES by about 55%.
@@ II.2 Interpretability of CGXAS in XANES Prediction
XANES is determined by both the absorption atom and its local atomic environment[RN230]. To analyze the interpretability of CGXAS on XANES prediction, it is essential to verify whether the features of absorption atom and local atomic environment can be identified from the latent site vectors after the information gathering by graph neural network. Here, the t-distributed stochastic neighbor embedding (t-SNE) method[RN323] is applied to reduce the high dimension latent site vectors into 2 t-SNE features, so that the they can be plotted in 2D space. Fig.3(b) shows the distribution of latent site vectors obtained from different absorption sites of various materials in the space of t-SNE features. Here, each element is marked by one specific color, which can be referred by its atomic number in the color bar. In the figure, the points with the same color tend to form a cluster, and clusters of different colors are separated obviously. It means that the information of absorption element is still kept in the latent site vector, which can be extracted by the t-SNE method.
@@ II.3 Transfer learning of CGXAS_Uni for Experiment-calibrated XANES Prediction
To assess the performance of transfer learning using a small experimental dataset, the dataset is randomly separated into training, validation and testing sets, with a fixed ratio of 6:1:1, for 20 times; and each separation is used to fine-tune and test an independent model. The average prediction losses in training, validation, and testing sets before and after transfer learning over the 20 models are shown in Fig.5(a), separately for CGXAS_Exp_S, CGXAS_Exp_Ti and CGXAS_Exp_Fe, where the error bar represents the standard deviation of loss. For CGXAS_Exp_S, the prediction losses before transfer learning are around 0.7, indicating that the simulated XANES has a large systematic error compared with the experiment measurement. The standard deviation of loss before transfer learning in the validation set (0.317) and testing set (0.305) are much larger than that in training set (0.0684), which is due to the smaller sample numbers in these two sets. After the transfer learning, the average prediction losses are reduced by about 90% to around 0.07, showing the good performance of CGXAS_Exp_S. For CGXAS_Exp_Ti and CGXAS_Exp_Fe, the prediction losses before transfer learning are around 0.05 and 0.07 respectively. These losses are much smaller than that of CGXAS_S_Exp, because the peak intensity of Ti and Fe K edge XANES is not as high as S K edge XANES in general, making the misalignment of peak yield smaller value of RSE than S K edge XANES. After transfer learning, the prediction losses are still reduced to around 0.03 for both Ti and Fe K edge XANES, showing an accordant trend with CGXAS_Exp_S.
Energy misalignment in the simulated XANES will highly influence the accuracy of some key features like absorption edge energy. To assess the ability of transfer learning to calibrate the energy scale, the absolute errors of predicted edge energy are also analyzed and shown in Fig.5(b). Before transfer learning, the average errors of edge energy for the XANES of these three elements are around 4 eV. After the transfer learning, the average errors of edge energy are largely reduced by about 55%, from around 4 eV to around 1.8 eV. The results show that the transfer learning can effectively reduce the energy misalignment for different containing materials. Although an average error of 1.8 eV is still not enough for practical application, these results show the feasibility of the proposed method to calibrate the XANES prediction without a very huge experiment dataset.
@@
To further illustrate the effect of transfer learning on experimental-calibrated XANES prediction, Fig.5(c)-(h) show the XANES spectra predicted by CGXAS_Exp_S, CGXAS_Exp_Ti and CGXAS_Exp_Fe with different RSE values, as well as the corresponded spectra predicted by CGXAS_Uni and experimental spectra. Before the transfer learning, most of the predicted XANES spectrum shows a similar shape as the experimental spectrum, but matches poorly in the absorption edge and peak positions. After the transfer learning even on a small experimental dataset, the absorption edge, the main peak positions and intensity are changed to achieve smaller difference with the experimental spectrum. For those examples with low RSE, the absorption edge and the main peak are corrected and fit well with the experimental spectrum. Some spectra have relatively high RSE but their edge and peak positions are still well corrected, where the high RSE mainly come from the large difference in main peak intensity. In general, for all the three cases of S, Ti, and Fe K edge XANES, the predicted XANES achieve better accordance with experimental results after transfer learning with a small experimental dataset, showing the good feasibility of our proposed method.
@@
Figure 5: Performance of the CGXAS model on the prediction of experimental XANES after transfer learning. (a) Prediction loss of S, Ti, and Fe K edge XANES (measured by RSE) before and after transfer learning in training set, validation set, and testing set. (b) Error of edge energy compared with experimental XANES before and after transfer learning in training set, validation set, and testing set. Here, the column height and error bar represent the average value and standard deviation of the prediction losses or energy errors calculated over 20 models trained with different data separations respectively. (a) and (b) share the same legend. (c)-(d) Examples of S K edge XANES predicted by CGXAS_Exp_S for the samples in testing sets with different RSE values, together with the spectrum predicted by CGXAS_Uni and the ground truth experimental XANES. (e)-(f) Examples of Ti K edge XANES predicted by CGXAS_Exp_Ti for the samples in testing sets. (g)-(h) Examples of Fe K edge XANES predicted by CGXAS_Exp_Fe for the samples in testing sets.

### v2

@@ Abstract
Theoretical simulation is helpful for accurate interpretation of experimental X-ray absorption near-edge structure (XANES) spectra that contain rich atomic and electronic structure information of materials. However, current simulation methods are usually too complex to give the needed accuracy and timeliness when a large amount of data need to be analyzed, such as for in-situ characterization of battery materials. To address these problems, artificial intelligence (AI) models have been developed for XANES prediction. However, instead of using experimental XANES data, the existing models are trained using simulated data, resulting in significant discrepancies between the predicted and experimental spectra. Also, the universality across different elements has not been well studied for such models. In this work, we firstly establish a crystal graph neural network, pre-trained on simulated XANES data covering 48 elements, to achieve universal XANES prediction with a low average relative square error of 0.020223; and then utilize transfer learning to calibrate the model using a small experimental XANES dataset. After calibration, the edge energy misalignment error of the predicted S, Ti and Fe K edge XANES is significantly reduced by about 80%. The method demonstrated in this work opens up a new way to achieve fast, universal, and experiment-calibrated XANES prediction.
@@ I Introduction
X-ray absorption spectroscopy (XAS) is a powerful tool to detect the electronic structures and atomic geometry of materials, including atom valence, bond length and coordination number[24]. It has the advantages of non-destructive, element selectivity, and high sensitivity[23]. In addition, XAS is available for various sample types, such as bulk materials, powders, and solutions[7]. Therefore, it has been widely applied to study the problems in different areas, for example, low dimensional materials characterization[48, 49], catalytic mechanism analysis[5, 53], relics identification[1], and biomedical analyses[38]. Especially, these unique advantages enable XAS to combine with the operando/in-situ techniques for real-time characterization of some complex systems, like electrochemistry reactions in batteries[54, 21, 42, 8]. According to the energy region with respect of the absorption edge, XAS can be divided into X-ray absorption near-edge structure (XANES) and extended X-ray absorption spectroscopy (EXAFS). EXAFS is mainly dominated by single scattering events, where coordination numbers and interatomic distances can be derived from Fourier transform for some materials with periodic structure[16, 46]. XANES is related to strong scattering processes as well as local atomic resonances, which is more complicated than EXAFS and hard to directly transformed to structure parameters[43]. Commonly, interpretation of experimental XANES is based on some empirical features, such as comparing edge shifts and peak positions with referenced spectra of standard materials[16].
@@
With the development of XAS theory and the corresponding calculation software like FEFF9[44], FDMNES[Bunău_2009], and OCEAN[50], the XANES can be quantitively simulated from a given crystal or molecule structure and taken as the referenced spectra in analysis. It helps to gain deeper insight into the relationship between local environment of absorbing atoms with spectra features. However, such calculation is time-consuming when the materials structure is complex. And the calculated spectra are sensitive to the selected approximation methods and calculation parameters, where a tedious parameters optimization process is needed to obtain reasonable results. In addition, XANES simulation faces challenges in aligning absolute energy scales and accounting for experimental broadening effects, resulting in inevitable discrepancies compared with experiments[34]. Without careful experimental calibration, the simulated XANES may yield false edge and peak features, misleading the XANES analysis. These shortcomings of current simulation methods have brought significant challenges for efficient and accurate XANES interpretation. This especially limits the in-situ XANES study of some essential but complex systems requiring massive data analysis, such as battery materials that experience constant structure changes during charge and discharge processes.
@@
The development of artificial intelligent (AI) has opened up a new paradigm to solve the similar complex problems in different areas, such as AlphaFold for protein structure prediction[22], GNoME for material stable structures discovery[35], and DeepH for Hamiltonian calculation in DFT[51]. Researchers have also made some attempts to apply AI methods in XANES prediction to avoid the complex calculation process. These developed models cover the predictions on C K edge XANES of amorphous carbon and organic molecules[28, 29], O and N K edge XANES of molecules[6], and 3d transition metals K edge XANES of transition metal compounds[25, 39, 40]. Although those models offer a more efficient and convenient way to obtain XANES, their universality is challenged for they only focus on limited range of absorption elements. The universality of AI based XANES prediction on other common elements need to be further studied. In addition, due to the shortage of experimental spectra, the existing models are trained based on the high throughput simulated XANES, where the spectra features and energy scale are not accurately calibrated. The uncorrected discrepancies between the training data and experiment measurement challenge the reliability of existing XANES prediction models. Therefore, the universality and fidelity of AI based XANES prediction should be improved before it can be put into practical use.
@@
In this work, we explored the universal and experiment-calibrated prediction of XANES through crystal graph neural network and transfer learning strategy. A crystal graph neural network model for XANES prediction (CGXAS) is firstly established, and the universality of CGXAS on common detected elements of K edge XANES is studied. Then, transfer learning strategy is adopted to calibrate the predicted XANES using a small experimental dataset. The framework of CGXAS and the training flowchart is represented in Fig.1. In CGXAS, the graph neural network (GNN) encodes the local environment information into a latent site vector for each atom in the crystal graph. Then, the multilayer perceptron (MLP) decodes the site vector into the corresponded K edge XANES of each site. The successive mask pooling layer is used to average the XANES at each site according to the element, enabling XANES prediction for different elements in one network. A universal dataset with 341405 simulated XANES covering 48 elements[33] is used for CGXAS training and evaluation. This study compares three distinct approaches for achieving universal XANES prediction: (1) Training a single model (CGXAS_Uni) on the full dataset for universal XANES prediction; (2) Training individual element-specific CGXAS models using the element-specific subsets and assembling them for universal XANES prediction, where model for element X is denoted as CGXAS_Spec_X; (3) Fine-tuning CGXAS_Uni model on each element-specific subset, leading to CGXAS_FT_X model that represents the fine-tuned version for element X. Among the three models, the single CGXAS_Uni model demonstrated the best overall accuracy and universality. To study the interpretability of CGXAS, the t-SNE method is utilized to analyze the latent site vectors. Furthermore, to bridge the gap between simulation and experiment, transfer learning is also adopted to adapt the pretrained CGXAS_Uni model with a small experimental XANES dataset of element X, where the model after transfer learning is denoted as CGXAS_Exp_X. Compared with traditional learning process, transfer learning can transfer the knowledge learned in other domains to enhance the prediction performance on target domain[37]. Since the prediction performance of neural network is closely related with training data size, transferring simulated XANES knowledge to experimental XANES prediction can compensate the insufficiency in experimental data quantity. Here, we demonstrate the ability of transfer learning to calibrate the XANES prediction of S, Ti and Fe K edge XANES using only 48, 40 and 45 experiment spectra for each element, where S, Ti, and Fe are the critical elements in the study of energy and catalysis. The calibrated CGXAS_Exp_S, CGXAS_Exp_Ti, and CGXAS_Exp_Fe models reduce the edge energy misalignment error of the predicted XANES by about 80%.
@@ II.2 Interpretability of CGXAS in XANES Prediction
XANES is determined by both the absorption atom and its local atomic environment[43]. To analyze the interpretability of CGXAS on XANES prediction, it is essential to verify whether the features of absorption atom and local atomic environment can be identified from the latent site vectors after the information gathering by graph neural network. Here, the t-distributed stochastic neighbor embedding (t-SNE) method[47] is applied to reduce the high dimension latent site vectors into 2 t-SNE features, so that the they can be plotted in 2D space. Fig.3(b) shows the distribution of latent site vectors obtained from different absorption sites of various materials in the space of t-SNE features. Here, each element is marked by one specific color, which can be referred by its atomic number in the color bar. In the figure, the points with the same color tend to form a cluster, and clusters of different colors are separated obviously. It means that the information of absorption element is still kept in the latent site vector, which can be extracted by the t-SNE method.
@@ II.3 Transfer learning of CGXAS_Uni for Experiment-calibrated XANES Prediction
To assess the performance of transfer learning using a small experimental dataset, the dataset is randomly separated into training, validation and testing sets, with a fixed ratio of 6:1:1, for 20 times; and each separation is used to fine-tune and test an independent model. The average value of metrics over the 20 independent models on each data split is used to evaluate the performance of transfer learning. Energy misalignment in the simulated XANES will highly influence the accuracy of some key features like absorption edge energy. To assess the ability of transfer learning to calibrate the energy scale, the absolute errors of predicted edge energy are firstly analyzed. By directly transferring the CGXAS_Uni model on the small experimental dataset of S K edge XANES, yielding the CGXAS_Exp_S model, the edge energy error in testing set is reduced from 3.8 eV to 1.4 eV. The transferred model is then applied to calibrate the simulated dataset of S K edge XANES. The CGXAS_Uni model is again finetuned on the calibrated simulated dataset and then further transferred on the experimental dataset to a new CGXAS_Exp_S model, resulting in a smaller edge energy error of 0.9 eV. The procedure of simulation data calibration and the multistep transfer learning can be repeated with the new CGXAS_Exp_S model, forming a closed cycle. Here, this procedure is cycled for 10 times, and the edge energy error is convergent to around 0.9 eV after 3 cycles. The CGXAS_Exp_S models from the cycle 6 with the lowest edge energy error is selected for further analysis in this work. The experiment-calibrated models for Ti (CGXAS_Exp_Ti) and Fe (CGXAS_Exp_Fe) are also trained in the same way. The edge energy error of CGXAS_Exp_S, CGXAS_Exp_Ti, and CGXAS_Exp_Fe in training, validation, and testing sets are shown in Fig.5(a). Before transfer learning, the average errors of edge energy for the XANES of these three elements are around 4 eV. The standard deviation of error before transfer learning in the validation set and testing set are much larger than that in training set, which is due to the smaller sample numbers in these two sets. After the transfer learning, the average errors of edge energy are largely reduced to around 0.8 eV by 80% for S K edge XANES. For Ti and Fe, the errors are reduced to around 1.1 eV. The standard deviation of errors is also reduced after transfer learning. The results show that the transfer learning can effectively reduce the energy misalignment for different containing materials.
@@
The average prediction losses in training, validation, and testing sets before and after transfer learning over the 20 models are shown in Fig.5(b), separately for CGXAS_Exp_S, CGXAS_Exp_Ti and CGXAS_Exp_Fe, where the error bar represents the standard deviation of loss. For CGXAS_Exp_S, the prediction losses before transfer learning are around 0.7, indicating that the simulated XANES has a large systematic error compared with the experiment measurement. The edge and peak of many predicted spectra severely misalign with experimental S K edge XANES spectra, especially for those with absorption edge appearing around 2480 eV. After the transfer learning, the average prediction losses are reduced by about 90% to around 0.07, showing the good performance of CGXAS_Exp_S. For CGXAS_Exp_Ti and CGXAS_Exp_Fe, the prediction losses before transfer learning are around 0.05 and 0.07 respectively. These losses are much smaller than that of CGXAS_S_Exp, because the peak intensity of Ti and Fe K edge XANES is not as high as S K edge XANES in general, making the misalignment of peak yield smaller value of RSE than S K edge XANES. After transfer learning, the prediction losses are still reduced to around 0.02 for both Ti and Fe K edge XANES, showing an accordant trend with CGXAS_Exp_S.
Although the experimental datasets have a wide distribution in the simulated spectrum space, there are still some regions without experimental samples. Without available experimental spectrum, it’s hard to directly evaluate the calibration performance of transfer learning in these regions. In order to estimate the generalization ability of our proposed method, we collect the prediction loss of each sample from testing sets in all 20 splits, to see whether the similarity of testing sample with the corresponded training and validation sets will influence the calibration performance. Here, the similarity is measured by the minimum distance of the testing sample to all the training and validation samples (DTV) in the simulated spectrum space (Fig.4), which is defined in Eq.3 in Method section. A smaller DTV value means that the testing sample has larger similarity with the training and validation sets. By this analysis, the performance of the calibrated models in the regions closed and far from the regions with experimental training data can be estimated, which can be used to indirectly assess the generalization ability of the model. The prediction losses measured by RSE of the test samples together with their DTV are shown in Fig.5(c) for the dataset of S. The blue and green histograms show the general distribution of DTV and prediction loss of these samples with dash line marking the average value. The DTV value in all the three cases have a wide distribution from 0 to 30. For the testing sets of S K edge XANES, comparing the samples of low (0-10), medium (10-20), and high (20-30) DTV values, their prediction losses all presents a similar distribution as the general distribution. The results show that the models have consistent performance distribution for samples with different similarities compared with training samples, implying a good generalization ability.
@@
To further illustrate the effect of transfer learning on experimental-calibrated XANES prediction, Fig.5(d)-(l) show the XANES spectra predicted by CGXAS_Exp_S, CGXAS_Exp_Ti and CGXAS_Exp_Fe with different RSE values, as well as the corresponded spectra predicted by CGXAS_Uni and experimental spectra. Before the transfer learning, most of the predicted XANES spectrum shows a similar shape as the experimental spectrum, but matches poorly in the absorption edge and peak positions. After the transfer learning even on a small experimental dataset, the absorption edge, the main peak positions and intensity are changed to achieve smaller difference with the experimental spectrum. For those examples with low RSE, the absorption edge and the main peak are corrected and fit well with the experimental spectrum (Fig.5(d), (g) and (j)). For those with medium RSE (around the average RSE), the edge position and the major shape of predicted spectra can coincide with the experimental spectra in general (Fig.5(e), (h) and (k)). The errors mainly come from the slight misalignment of peaks, missing of some fine structures or peak intensity difference. From Fig.5(c), the majority of predicted spectra are in the region of low and medium RSE. Some spectra have relatively high RSE but their edge and peak positions are still well corrected (Fig.5(f) and (i)), where the high RSE mainly come from the large difference in main peak intensity. For some pre-edge structures or post edge peaks with lower intensity, the calibrated models may still yield large error, which requires further optimization in future study. In general, for all the three cases of S, Ti, and Fe K edge XANES, the predicted XANES achieve better accordance with experimental results after transfer learning with a small experimental dataset, showing the good feasibility of our proposed method.
@@
Figure 5: Performance of the CGXAS model on the prediction of experimental XANES after transfer learning. (a) Error of edge energy compared with experimental XANES before and after transfer learning in training set, validation set, and testing set. Here, the column height and error bar represent the average value and standard deviation of the prediction losses or energy errors calculated over 20 models trained with different data separations respectively. (a) and (b) share the same legend. (b) Prediction loss of S, Ti, and Fe K edge XANES (measured by RSE) before and after transfer learning in training set, validation set, and testing set. (c) Scatter plot of the prediction loss for different samples in S K edge XANES testing sets from 20 random splits and their distance to the training and validation set (DTV). The histograms on the top and right show the distribution of these on the variances of the x axis and y axis, respectively. And the dash lines mark the average values of the variables in x and y axis. (d)-(f) Examples of S K edge XANES predicted by CGXAS_Exp_S for the samples in testing sets with different RSE values, together with the spectrum predicted by CGXAS_Uni and the ground truth experimental XANES. (g)-(i) Examples of Ti K edge XANES predicted by CGXAS_Exp_Ti for the samples in testing sets. (j)-(l) Examples of Fe K edge XANES predicted by CGXAS_Exp_Fe for the samples in testing sets.